# UVIP: Model-Free Approach to Evaluate Reinforcement Learning Algorithms

### v4

@@ A.5 The covering radius of randomly distributed points over a cube
The following proposition is a particular case of the result (Reznikov & Saff, 2016, Theorem 2.1). We repeat the arguments from that paper and give explicit expressions for the constants.
Proposition A.3.
@@
Proof.
Let $\mathcal{E}_{n}=\mathcal{E}_{n}(\mathsf{X})$ be a maximal set of points such that for any $y,z\in\mathcal{E}_{n}$ we have $|y-z|\geq 1/n$. Then for any $x\in\mathsf{X}$ there exists a point $y\in\mathcal{E}_{n}$ such that $|x-y|\leq 1/n$. Denote by $B(x,r)$ a ball centred at $x\in\mathsf{X}$ of radius $r$ (w.r.t. ||$\cdot$|| ⋅ |) and
$\Phi(r)=\frac{r^{d_{\mathsf{X}}}\pi^{{d_{\mathsf{X}}}/2}}{2^{d_{\mathsf{X}}}\Gamma({d_{\mathsf{X}}}/2+1)},\,r\in[0,\infty)$
@@
Suppose that $\rho(\mathsf{X}_{N},\mathsf{X})>2/n$. Then there exists a point $y\in\mathsf{X}$ such that $\mathsf{X}_{N}\cap B(y,2/n)=\emptyset$. Choose a point $x\in\mathcal{E}_{n}$ such that $|x-y|<1/n$. Then $B(x,1/n)\subset B(y,2/n)$, and so the ball $B(x,1/n)$ doesn’t intersect $\mathsf{X}_{N}$. Hence, $\mathsf{X}_{N}\cap B(x,1/(3n))=\emptyset$. Therefore,
@@
$\Phi^{-1}(r)=\frac{2}{\sqrt{\pi}}\Gamma^{1/{d_{\mathsf{X}}}}({d_{\mathsf{X}}}/2+1)r^{1/{d_{\mathsf{X}}}}\leq 2\sqrt{\mathrm{e}{d_{\mathsf{X}}}/\pi}(\mathrm{e}r)^{1/{d_{\mathsf{X}}}}$
@@
It remains to take $\alpha=1+p/{d_{\mathsf{X}}}$ to obtain bound
@@
Hence, (33) follows. To prove (34) it remains to apply Markov’s inequality. ∎
@@ A.6 Auxiliary results
Proof.
Consider the case $\Delta<1$. In this case
@@
$\lesssim$ $\Delta^{50}+(\Delta-\Delta^{100/2})\sqrt{\log(1+2/\Delta^{100})}$
$\lesssim$ $\Delta\sqrt{\log(1+1/\Delta)}.$
If $\Delta>1$,
$\int_{0}^{\Delta}\sqrt{\log(1+1/x)}\mathrm{d}x$ $=$ $\int_{0}^{1}\sqrt{\log(1+1/x)}\mathrm{d}x+\int_{1}^{\Delta}\sqrt{\log(1+1/x)}\mathrm{d}x$
$\lesssim$ $2+(\Delta-1)\sqrt{\log 2}\lesssim\Delta.$
@@ B.1 Environments description
Garnet example is an MDP with randomly generated transition probability kernel ${\rm P}^{a}$ with finite $\mathsf{X}$ state space and $\mathsf{A}$ action space. This example is described with a tuple $\left\langle N_{S},N_{A},N_{B}\right\rangle$. The first two parameters specify the number of states and actions respectively. The last parameter is responsible for the number of states an agent can go to from state $x\in\mathsf{X}$ by performing action $a\in\mathsf{A}$. In our case we used $N_{S}=20,N_{A}=5,N_{B}=2,\gamma=0.9$. The reward matrix $r^{a}(x)$ is set according to the following principle: first, for all state-action pairs, the reward is set uniformly distributed on [0, 1]. Then the pairs are randomly chosen, for which the reward will be increased several times.
@@
The agent moves in a grid world, where some squares of the lake are walkable, but others lead to the agent falling into the water, so the game restarts. Additionally, the ice is slippery, so the movement direction of the agent is uncertain and only partially depends on the chosen direction. The agent receives 10 points only for finding a path to a goal square, for falling into the hole it doesn’t receive anything. We used the built-in $4\times 4$ map and 4 actions for the agent to perform on each state if available (right, left, up and down). For this experiment we assume that the reward matrix $r^{a}(x)$ is known, $\gamma$-factor is set to be 0.9.
@@
Chain is a finite MDP where agent can move only to 2 adjacent states performing 2 actions from each state (right and left). Every chain has two terminal states at the ends. For transition to the terminal states agent receives 10 points and episode ends, otherwise the reward is equal to +1. Also, there is pnoise in the system, that is the agent performs a uniformly-random action with probability p. For experiments with chains we set $\gamma$-factor to 0.8, to be sure that Pichard iterations converge.
@@
NRoom is a discrete Grid-World environment with connected rooms and with one large reward in a single room and small rewards elsewhere. Also, there are traps which lead to the terminal states. The example of this environment is illustrated by Figure 4. At each state there are four actions to turn on left, right, up and down. With a small probability the chosen action is ignored and uniformly random action is chosen.
Figure 4: Discrete Grid-World environment with connected rooms. The red squares are traps which are terminal states.
@@
CartPole is an example of the environment with a finite action space and infinitely large state space. In this environment agent can push cart with pole on it to the left or right direction and the target is to hold the pole up as long as possible. Reward equals to 1 is gain every time until failing or the end of episode. In fact, CartPole hasn’t any specific stochastic dynamic, because transitions are deterministic according to actions, so for non degenerate case we should add some noise and we apply normally distributed random variable to the angle. LD(Linear Deterministic) policy can be expressed as I$\{3\cdot\theta+\dot{\theta}>0\}$, where $\theta$ is an angle between pole and normal to cart.
@@
The environment consists of two joints, or two links. The torque is applied to the binding between the joints. The action space is discrete and consists of three kinds of torques: left, right and none. The state space is 6 dimensional, representing two angles (sine and cosine) characterizing the links position and the angles’ velocities. Each episode starts with the small perturbations of the parameters near the resting state having both of the joints in a downward position. The agents’ goal is to reach a given boundary from above in a minimum amount of time with its lower point of the link. At each timestep a robot has a reward equal to -1, and it gets 0 in a terminal state, when the boundary has been reached. Also to make the environment stochastic, a random uniform torque from $-1$ to $1$ is added to the force at each step.
@@
TwinRooms is Grid-World environment with continuous state space. It is composed of two rooms separated by a wall, such that $\mathsf{X}=([0,1-\Delta]\cup[1+\Delta,2])\times[0,1]$ where $0.12\Delta=$ 0.1 is the width of the wall, as illustrated by Figure 5. There are four actions: left, right, up, and down, each one resulting to a displacement of 0.1 in the corresponding direction. A two-dimensional Gaussian noise is added to the transitions, and, in each room, there is a single region with non-zero reward. The agent has 0.5 probability of starting in each of the rooms, and the starting position is at the room’s bottom left corner.
@@ B.2 Experimental setup
Environment $M_{1}$ $M_{2}$ discount $\gamma$ $N$
@@ B.3 Auxiliary algorithms
For the sake of completeness we provide here the Value iteration algorithm (Szepesvári (2010), Chapter 1), used in Section 6 for tabular environments.
Input: ${\rm P}^{a}$, $r$, $\varepsilon$
Result: $V^{\star}$
Initialize: $\forall x\in\mathsf{X},\ V^{\star}_{0}(x)=0$
$k=1$
while $\left\|V^{\star}_{k}-V^{\star}_{k-1}\right\|>\varepsilon$ do
       for $x\in\mathsf{X},\ a\in\mathsf{A}$ do
             $Q(x,a)=r^{a}(x)+\gamma\sum\limits_{y\in\mathsf{X}}{\rm P}^{a}(y|x)V^{\star}_{k-1}(y)$
             $V^{\star}_{k}(x)=\max\limits_{a\in\mathsf{A}}{Q(x,a)}$
       end for
      k = k + 1
end while
Algorithm 3 Value Iteration algorithm

### v5

@@ A.5 The covering radius of randomly distributed points over a cube
The following proposition is a particular case of the result (Reznikov and Saff, 2015, Theorem 2.1). We repeat the arguments from that paper and give explicit expressions for the constants.
Proposition A.3
@@
Proof
Let $\mathcal{E}_{n}=\mathcal{E}_{n}(\mathsf{X})$ be a maximal set of points such that for any $y,z\in\mathcal{E}_{n}$ we have $|y-z|\geq 1/n$. Then for any $x\in\mathsf{X}$ there exists a point $y\in\mathcal{E}_{n}$ such that $|x-y|\leq 1/n$. Denote by $B(x,r)$ a ball centred at $x\in\mathsf{X}$ of radius $r$ (w.r.t. $|\cdot|$) and
$\Phi(r)=\frac{r^{d_{\mathsf{X}}}\pi^{{d_{\mathsf{X}}}/2}}{2^{d_{\mathsf{X}}}\Gamma({d_{\mathsf{X}}}/2+1)},\,r\in[0,\infty).$
@@
Suppose that $\rho(\mathsf{X}_{N},\mathsf{X})>2/n$. Then there exists a point $y\in\mathsf{X}$ such that $\mathsf{X}_{N}\cap B(y,2/n)=\emptyset$. Choose a point $x\in\mathcal{E}_{n}$ such that |xy|<1/n|x-y|<1/n. Then $B(x,1/n)\subset B(y,2/n)$, and so the ball $B(x,1/n)$ doesn’t intersect $\mathsf{X}_{N}$. Hence, $\mathsf{X}_{N}\cap B(x,1/(3n))=\emptyset$. Therefore,
@@
$\Phi^{-1}(r)=\frac{2}{\sqrt{\pi}}\Gamma^{1/{d_{\mathsf{X}}}}({d_{\mathsf{X}}}/2+1)r^{1/{d_{\mathsf{X}}}}\leq 2\sqrt{\mathrm{e}{d_{\mathsf{X}}}/\pi}(\mathrm{e}r)^{1/{d_{\mathsf{X}}}},$
@@
It remains to take $\alpha=1+p/{d_{\mathsf{X}}}$ to obtain the bound
@@
Hence, (33) follows. To prove (34) it remains to apply Markov’s inequality.
@@ A.6 Auxiliary results
Proof
Consider first the case $\Delta<1$. In this case
@@
$\lesssim$ $\Delta^{50}+(\Delta-\Delta^{100/2})\sqrt{\log(1+2/\Delta^{100})}\lesssim\Delta\sqrt{\log(1+1/\Delta)}.$
Second, if $\Delta>1$,
$\int_{0}^{\Delta}\sqrt{\log(1+1/x)}\mathrm{d}x=\int_{0}^{1}\sqrt{\log(1+1/x)}\mathrm{d}x+\int_{1}^{\Delta}\sqrt{\log(1+1/x)}\mathrm{d}x\lesssim\Delta\,.$
@@ B.1 Environment description
The Garnet example is an MDP with randomly generated transition probability kernel ${\rm P}^{a}$ with finite state space $\mathsf{X}$ and action space $\mathsf{A}$. This example is described by a tuple $\left\langle N_{S},N_{A},N_{B}\right\rangle$. The first two parameters specify the number of states and actions, respectively. The last parameter is responsible for the number of states to which an agent can go from state $x\in\mathsf{X}$ by performing action $a\in\mathsf{A}$. In our case, we used $N_{S}=20,N_{A}=5,N_{B}=2,\gamma=0.9$. The reward matrix $r^{a}(x)$ is set according to the following principle: for all state-action pairs, the reward is set to be uniformly distributed on [0, 1].
@@
The agent moves in a grid world, where some squares of the lake are walkable, but others lead to the agent falling into the water, so the game restarts. Additionally, the ice is slippery, so the movement direction of the agent is uncertain and only partially depends on the chosen direction. The agent receives 10 points only for finding a path to a goal square, whereas for falling into a hole it does not receive anything. We used the built-in $4\times 4$ map and 4 actions for the agent to perform in each state, if available (right, left, up and down). For this experiment, we assume that the reward matrix $r^{a}(x)$ is known, and the $\gamma$-factor is set to be 0.9.
@@
Chain is a finite MDP where the agent can move only to 2 adjacent states, performing 2 actions from each state (right and left). Every chain has two terminal states at the ends. For a transition to the terminal states, the agent receives 10 points and the episode ends, otherwise the reward is equal to +1. Also, there is $p\%$ noise in the system, that is, the agent performs a uniformly random action with probability $p$. For experiments with chains, we set the $\gamma$-factor to 0.8, to ensure that Picard iterations converge.
@@
NRoom is a discrete grid-world environment with connected rooms and with one large reward in a single room and small rewards elsewhere. Also, there are traps which lead to terminal states. At each state there are four actions: left, right, up, and down. With a small probability, the chosen action is ignored and a uniformly random action is chosen.
@@
[…]
@@
The environment consists of two joints, or two links. The torque is applied to the binding between the joints. The state space is six-dimensional, representing two angles (sine and cosine) characterizing the links’ positions and the angles’ velocities. Each episode starts with small perturbations of the parameters near the resting state with both of the joints in a downward position. At each time step the robot has a reward equal to -1, and it gets 0 in a terminal state, when the boundary has been reached. Also, to make the environment stochastic, a random uniform torque from 1-1 to $1$ is added to the force at each step.
@@
TwinRooms is a grid-world environment with continuous state space. It is composed of two rooms separated by a wall, such that $\mathsf{X}=([0,1-\Delta]\cup[1+\Delta,2])\times[0,1]$ where $2\Delta=0.1$ is the width of the wall, as illustrated by Figure 4. There are four actions: left, right, up, and down, each one resulting in a displacement of 0.1 in the corresponding direction. A two-dimensional Gaussian noise is added to the transitions, and, in each room, there is a single region with non-zero reward. The agent has 0.5 probability of starting in each of the rooms, and the starting position is at the room’s bottom-left corner.
@@ B.2 Experimental setup
Environment $M_{1}$ $M_{2}$ discount $\gamma$ NN
@@ B.3 Auxiliary algorithms
In Section 7 we use the Value Iteration algorithm from Szepesvári (2010), Chapter 1.